OpenStudy (anonymous):

f(x)=px^2+qx-5. f(-1) and f(1)=-9 (a)find the values of p and q (b)the values of x for which f(x)=1.

OpenStudy (whpalmer4):

f(-1) = -9 = p(-1)^2 + q(-1) - 5 = p - q - 5 f(1) = -9 = p(1)^2 + q(1) - 5 = p + q - 5 therefore p - q - 5 = p + q - 5 = -9 p - q - 5 = -9 p + q - 5 = -9 ------------- 2p - 10 = -18 2p = -8 p = -4 -4 -q -5 = -9 q = 0 Checking our work, -4(-1)^2 + 0(-1) - 5 = -9 for part b, we need to set f(x) = 1 = px^2 + qx -5 1 = -4x^2 + 0x - 5 0 = -4x^2 + 0x - 6 use the quadratic formula to solve for x